OpenStudy (anonymous):

Solve this equation by using the quadratic form. x^4 - 144 = 0

OpenStudy (whpalmer4):

Substitute \[u=x^2\] and solve via quadratic formula, then undo substitution and solve again

OpenStudy (anonymous):

Hint: Start off by factoring the equation through "Difference of Squares":\[\bf x^4-144=0 \implies x^4-12^2=0 \implies (x^2)^2-12^2=0\]

OpenStudy (anonymous):

Note that the difference of squares for \(a\) and \(b\) is: \(a^2-b^2=(a-b)(a+b)\). After factoring, equate each factor with 0 and solve for x.